# MultiGran-STGCNFog: Towards Accurate and High-Throughput Inference for Multi-Granular Spatiotemporal Traffic Forecasting

### v1

@@ 4.2.1 Multi-scale Traffic Data Modeling
$A(i,j)=\exp\left(-\frac{d_{ij}}{\text{var}(d_{i})}\right),\quad 0\leq A(i,j)\leq 1.$ (9)
@@
$L=D_{\text{deg}}-A,\quad L_{\text{norm}}=D_{\text{deg}}^{-\frac{1}{2}}LD_{\text{deg}}^{-\frac{1}{2}}.$ (10)
@@
$\min\sum_{i=1}^{n}\sum_{j=1}^{k}\left\|\textbf{H}_{i}-\mu_{j}\right\|^{2},$ (11)
where $\mu_{j}$ is the centroid of cluster $j$. Each cluster is treated as a node, $\mathbf{X}_{i}\in\mathbb{R}^{T\times d}$, in the coarser spatial scale, and features are aggregated through pooling operations:
$\mathbf{X}_{\text{min}}^{(j)}$ $=\min_{i\in\mathcal{C}_{j}}\mathbf{X}_{i},\;\mathbf{X}_{\text{mean}}^{(j)}=\frac{1}{|\mathcal{C}_{j}|}\sum_{i\in\mathcal{C}_{j}}\mathbf{X}_{i},\;\mathbf{X}_{\text{max}}^{(j)}=\max_{i\in\mathcal{C}_{j}}\mathbf{X}_{i},$ (12)
@@
4.2.2 Dynamic Graph Generation Block
Unlike the traditional graph convolution operator, defined by a static graph structure $\hat{A}$:
$H^{(l+1)}=\sigma\left(\hat{D}^{-\frac{1}{2}}\hat{A}\hat{D}^{-\frac{1}{2}}H^{(l)}W^{(l)}\right).$ (13)
@@ 4.2.2 Dynamic Graph Generation Block
$A_{ij}=\text{ReLU}(\mathbf{v}_{1,i}\cdot\mathbf{v}_{2,j})=\text{ReLU}\left(\sum_{k=1}^{d}v_{1,i,k}\cdot v_{2,j,k}\right).$ (14)
4.2.3 Feature Extraction Branches
@@
$X_{\text{temp}}=\tanh(W_{\text{time}}*X+b_{\text{time}}),$ (15)
@@
$X^{\prime}_{\text{temp}}=\tanh(X_{\text{temp}_{1}})\cdot\sigma(X_{\text{temp}_{2}}).$ (16)
@@
$X_{\text{gcn}}=\tilde{D}^{-1/2}A_{\text{dynamic}}\tilde{D}^{-1/2}X^{\prime}_{\text{temp}}W_{\text{gcn}}.$ (17)
@@
Having $A_{\text{dynamic}}$ representing weighted spatial connections between nodes, it is similar to the temporal attention mechanism that dynamically assigns weights to different time steps, leveraging Scaled Dot-Product Attention with query (Q), key (K), and value (V) matrices:
@@
$\begin{split}Q=X_{\text{gcn}}W_{Q},\quad K=X_{\text{gcn}}W_{K},\quad V=X_{\text{gcn}}W_{V},\\ \text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V.\end{split}$ (18)
@@
C-GF blocks integrate multi-granular feature representations through learnable gates that regulate the contribution of each input feature set to the final fused representation. Features from spatial scales: $X_{L1}$, $X_{L2}$, and $X_{L3}$ are first concatenated to form a joint representation $X_{\text{concat}}$. Subsequently, the gating mechanism adjusts the contribution by the learnable weight matrix $W_{g}$ followed by a fusion step:
$\mathbf{G}=\sigma(W_{g}\cdot X_{\text{concat}}+b_{g}),$ (19)
$X_{\text{fused}}=\mathbf{G}_{L1}\odot X_{L1}+\mathbf{G}_{L2}\odot X_{L2}+\mathbf{G}_{L3}\odot X_{L3},$ (20)
$X_{\text{fusion}}=\sigma(W_{\text{fusion}}\cdot X_{\text{fused}}+b_{\text{fusion}}).$ (21)
4.2.4 Traffic Forecasting Head
@@
In the forecasting head, to enable dynamical adjustment on the contribution of each time scale based on their importance, learnable weight matrices are exploited to compute the final prediction operating the output features $x_{\text{hour}}$, $x_{\text{day}}$, and $x_{\text{week}}$:
$x=W_{\text{hour}}\odot x_{\text{hour}}+W_{\text{day}}\odot x_{\text{day}}+W_{\text{week}}\odot x_{\text{week}}.$ (22)
4.3 Cross-device execution scheduling
@@
Explained by (8), $(O^{*},S^{*})$ is crucial in determining how layers should be orchestrated and assigned. It is worth noting that previous works often ignored the optimization for the layer execution order $O$. However, our experiments show that the layer execution order plays a decisive role regarding the pipeline inference throughput. We model the input-output dependencies between layers using a Directed Acyclic Graph (DAG), where each node represents a layer $l_{j}$. If the output of layer $l_{j}$ is required as input for layer $l_{j+1}$, a directed edge is established from $l_{j}$ to $l_{j+1}$. This DAG structure captures the sequential and parallel execution constraints between layers. As the number of layers $L$ increases, the search space for both $O^{*}$ and $S^{*}$ grows exponentially. Hence, we propose GA-DPHDS, an algorithm utilizing the NSGA-II genetic algorithm with dynamic programming for optimization.
@@
First, the population Pop is initialized randomly, where each individual $\text{ind}_{i}$ represents a layer execution order $O$, and individuals must satisfy the topological ordering constraints imposed by the DAG. For each $\text{ind}_{i}$, the overall inference time $T_{\text{overall}}(\text{ind}_{i})$, which is the longest device execution time; and the load balancing factor $\sigma(\text{ind}_{i})$, the standard deviation of execution times across all devices, are calculated:
$T_{\text{overall}}=\max_{d_{i}\in\mathcal{D}}\max\big{(}$ $\sum_{l_{j}\in\text{stage}_{i}}T_{\text{comp}}(l_{j},d_{i}),$ (23)
@@
$\sigma(\text{ind}_{i})=\sqrt{\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}\left(T_{\text{exec}}(d_{i})-\overline{T_{\text{exec}}}\right)^{2}}.$ (24)
@@
$\text{Distance}(\text{ind}_{i})=$ $\,\frac{T_{\text{overall}}(\text{ind}_{i+1})-T_{\text{overall}}(\text{ind}_{i-1})}{T_{\text{overall, max}}-T_{\text{overall, min}}}$ (25)
@@
For each individual $\text{ind}_{i}$, we leverage dynamic programming to further optimize the layer-device scheduling strategy $S$, defining a two-dimensional dynamic programming table $T$, where $T[i][j]$ represents the minimal maximum execution time when assigning the first $i$ layers to the first $j$ devices. The state transition recurrence is given by:
$T[i][j]=\min_{k<i}\left(\max\left(T[k][j-1],\right.\right.$ (26)
$\left.\left.\qquad\qquad\quad T_{\text{stage}}(k+1,i,D_{j})\right)\right).$
@@
1) The minimal maximum execution time of the first $k$ layers on first $j-1$ devices, $T[k][j-1]$.
@@
$T_{\text{stage}}(k+1,\ i,\ D_{j})=\max\left\{\begin{aligned} &T_{\text{comp}}(k+1,\ i,\ D_{j}),\\ &T_{\text{comm}}(D_{j-1},\ D_{j})\end{aligned}\right\}.$ (27)
@@
$T_{\text{comp}}(k+1,\ i,\ D_{j})=\sum_{l=k+1}^{i}T_{\text{comp}}(l,\ D_{j}),$ (28)
@@
$T_{\text{comm}}(D_{j-1},D_{j})=\begin{cases}\frac{\text{output}_{k}\cdot B_{\mu}}{b_{j-1,j}},&\text{if }D_{j-1}\neq D_{j},\\ 0,&\text{if }D_{j-1}=D_{j}.\end{cases}$ (29)
@@
5 Experiments
The experiments section aims to: (1) Validate the performance of MultiGran-STGCN. (2) Demonstrate the necessity of GA-DPHDS against homogeneous counterparts. (3) Explore the impact of device network bandwidth on overall inference throughput and identify system bottlenecks. (4) Evaluate the effectiveness of our optimization strategies.
Table 1: Heterogeneous device table

### v2

@@ 4.2.1. Multi-scale Traffic Data Modeling
(9) $A(i,j)=\exp\left(-\frac{d_{ij}}{\text{var}(d_{i})}\right),\quad 0\leq A(i,j)\leq 1.$
@@
(10) $L=D_{\text{deg}}-A,\quad L_{\text{norm}}=D_{\text{deg}}^{-\frac{1}{2}}LD_{\text{deg}}^{-\frac{1}{2}}.$
@@
(11) $\min\sum_{i=1}^{n}\sum_{j=1}^{k}\left\|\textbf{H}_{i}-\mu_{j}\right\|^{2},$
where $\mu_{j}$ is the centroid of cluster $j$. Each cluster is treated as a node, $\mathbf{X}_{i}\in\mathbb{R}^{T\times d}$, in the coarser spatial scale, and features are aggregated through a series of pooling operations:
(12) $\mathbf{X}_{\text{min}}^{(j)}$ $=\min_{i\in\mathcal{C}_{j}}\mathbf{X}_{i},\;\mathbf{X}_{\text{mean}}^{(j)}=\frac{1}{|\mathcal{C}_{j}|}\sum_{i\in\mathcal{C}_{j}}\mathbf{X}_{i},\;\mathbf{X}_{\text{max}}^{(j)}=\max_{i\in\mathcal{C}_{j}}\mathbf{X}_{i},$
@@
4.2.2. Dynamic Graph Generation Block
Unlike traditional graph convolution operators, defined by a static graph structure $\hat{A}$:
(13) $H^{(l+1)}=\sigma\left(\hat{D}^{-\frac{1}{2}}\hat{A}\hat{D}^{-\frac{1}{2}}H^{(l)}W^{(l)}\right).$
@@ 4.2.2. Dynamic Graph Generation Block
(14) $A_{ij}=\text{ReLU}(\mathbf{v}_{1,i}\cdot\mathbf{v}_{2,j})=\text{ReLU}\left(\sum_{k=1}^{d}v_{1,i,k}\cdot v_{2,j,k}\right).$
4.2.3. Feature Extraction Branches
@@
(15) $X_{\text{temp}}=\tanh(W_{\text{time}}*X+b_{\text{time}}),$
@@
(16) $X^{\prime}_{\text{temp}}=\tanh(X_{\text{temp}_{1}})\cdot\sigma(X_{\text{temp}_{2}}).$
@@
(17) $X_{\text{gcn}}=\tilde{D}^{-1/2}A_{\text{dynamic}}\tilde{D}^{-1/2}X^{\prime}_{\text{temp}}W_{\text{gcn}}.$
@@
Having $A_{\text{dynamic}}$ representing weighted spatial connections among nodes, it is similar to the temporal attention mechanism that dynamically assigns weights to different time steps by Scaled Dot-Product Attention with queries (Q), keys (K), and values (V):
@@
(18) $\begin{split}Q=X_{\text{gcn}}W_{Q},\quad K=X_{\text{gcn}}W_{K},\quad V=X_{\text{gcn}}W_{V},\\ \text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V.\end{split}$
@@
With the extraction of spatiotemporal dependencies, it becomes crucial to fuse multi-granular feature representations. C-GF blocks integrate multi-granular feature representations through learnable gates that regulate the contribution of each input feature set to the final fused representation. Features from spatial scales: $X_{L1}$, $X_{L2}$, and $X_{L3}$ are first concatenated to form a joint representation $X_{\text{concat}}$. Subsequently, the gating mechanism adjusts the contribution by the learnable weight matrix $W_{g}$ followed by a fusion step:
(19) $\mathbf{G}=\sigma(W_{g}\cdot X_{\text{concat}}+b_{g}),$
(20) $X_{\text{fused}}=\mathbf{G}_{L1}\odot X_{L1}+\mathbf{G}_{L2}\odot X_{L2}+\mathbf{G}_{L3}\odot X_{L3},$
(21) $X_{\text{fusion}}=\sigma(W_{\text{fusion}}\cdot X_{\text{fused}}+b_{\text{fusion}}).$
4.2.4. Traffic Forecasting Head
@@
In the forecasting head, to enable dynamical adjustment on the contribution of each time scale based on their importance, learnable weight matrices are exploited to compute the final prediction operating the output features $x_{\text{hour}}$, $x_{\text{day}}$, and $x_{\text{week}}$ from different feature extraction branches:
(22) $x=W_{\text{hour}}\odot x_{\text{hour}}+W_{\text{day}}\odot x_{\text{day}}+W_{\text{week}}\odot x_{\text{week}}.$
Figure 3. The illustrative diagram of inference acceleration through pipeline parallelism.
[…]
Scheduling figure
1
Input : $L$: set of layers, $D$: set of devices, $P$: processing time matrix, $B$: bandwidth between devices, $N_{\text{pop}}$: population size, $N_{\text{gen}}$: number of generations, $P_{c}$: crossover probability, $P_{m}$: mutation probability
Output : $S^{*}$: optimal scheduling strategy
2
3Initialize population Pop of size $N_{\text{pop}}$;
4 for each individual $\text{ind}_{i}$ in Pop do
5   Decode $\text{ind}_{i}$ to get layer execution order $O_{i}$;
6   Apply DP to get scheduling strategy $S_{i}$;
7   Calculate $T_{\text{overall}}(\text{ind}_{i})$ and $\sigma(\text{ind}_{i})$;
8
9for $\text{gen}=1$ to $N_{\text{gen}}$ do
10   Perform fast non-dominated sorting on Pop;
11   Calculate $\text{Distance}(ind_{i})$ for each $\text{ind}_{i}$;
12   Select parents based on rank and $\text{Distance}(ind_{i})$;
13   Apply crossover and mutation with $P_{c}$ and $P_{m}$;
14 for each offspring $\text{ind}^{\prime}_{i}$ do
15      Decode $\text{ind}^{\prime}_{i}$ to obtain new layer order $O^{\prime}_{i}$;
16      Apply DP to get scheduling strategy $S^{\prime}_{i}$;
17      Calculate $T_{\text{overall}}(\text{ind}^{\prime}_{i})$ and $\sigma(\text{ind}^{\prime}_{i})$;
18
19  Integrate population and offspring;
20   Select top $N_{\text{pop}}$ individuals for the next generation;
21
22return optimal $(O^{*},S^{*})$ based on $T_{\text{overall}}$ and $\sigma$;
23
24Dynamic Programming for Scheduling Strategy:
25 Initialize DP table $T[i][j]$ with infinity;
26 Set $T[0][0]\leftarrow 0$;
27 for $i=1$ to $|N|$ do
28 for $j=1$ to $|M|$ do
29    for $k=0$ to i1i-1 do
30       $T_{\text{comp}}\leftarrow\sum_{l=k+1}^{i}P[l][D_{j}]$;
31       $T_{\text{comm}}\leftarrow\frac{\text{output}_{k}\cdot B_{\mu}}{b_{j-1,j}}$;
32       $T_{\text{stage}}\leftarrow\max(T_{\text{comp}},\ T_{\text{comm}})$;
33         Update DP table $T[i][j]\leftarrow\min\left(T[i][j],\ \max\left(T[k][j-1],\ T_{\text{stage}}\right)\right)$;
34
35
36
37Backtrack $T[i][j]$ to find the optimal $S^{o}$ given a $O$;
38 return $S^{o}$ and min-max stage execution time;
Algorithm 1 GA-DPHDS: Genetic Algorithm with Dynamic Programming for Heterogeneous Device Scheduling
4.3. Cross-device execution scheduling
Tackling the prolonged inference time due to additional feature extraction branches, we accelerate the inference process of MultiGran-STGCN through pipeline parallelism, as shown in Fig. 3. Specifically, the entire traffic forecasting model is partitioned into atomic unit layers and allocated to heterogeneous fog devices as stages to perform distributed pipeline-parallel inference.
@@ 4.3. Cross-device execution scheduling
Explained by (8), $(O^{*},S^{*})$ is vital in determining how layers should be orchestrated and assigned. It is worth noting that previous works often ignored the optimization for the layer execution order $O$. However, our experiments show that the layer execution order plays a decisive role regarding the pipeline inference throughput. We model the input-output dependencies between layers using a Directed Acyclic Graph (DAG), where each node represents a layer $l_{j}$. If the output of layer $l_{j}$ is required as input for layer $l_{j+1}$, a directed edge is established from $l_{j}$ to $l_{j+1}$. This DAG structure captures the sequential and parallel execution constraints between layers. As the number of layers $L$ increases, the search space for both $O^{*}$ and $S^{*}$ grows exponentially. Hence, we propose GA-DPHDS, an algorithm utilizing the NSGA-II genetic algorithm with dynamic programming for optimization, depicted by Algorithm 1.
@@
First, the population Pop is randomly initialized, where each individual $\text{ind}_{i}$ represents a layer execution order $O$, and individuals must satisfy the topological ordering constraints imposed by the DAG. For each $\text{ind}_{i}$, the overall inference time $T_{\text{overall}}(\text{ind}_{i})$, longest device execution time; and the load balancing factor $\sigma(\text{ind}_{i})$, the standard deviation of execution times across devices, are calculated:
(23) $T_{\text{overall}}=\max_{d_{i}\in\mathcal{D}}\max\big{(}$ $\sum_{l_{j}\in\text{stage}_{i}}T_{\text{comp}}(l_{j},d_{i}),$
@@
(24) $\sigma(\text{ind}_{i})=\sqrt{\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}\left(T_{\text{exec}}(d_{i})-\overline{T_{\text{exec}}}\right)^{2}}.$
@@
(25) $\text{Distance}(\text{ind}_{i})=$ $\,\frac{T_{\text{overall}}(\text{ind}_{i+1})-T_{\text{overall}}(\text{ind}_{i-1})}{T_{\text{overall, max}}-T_{\text{overall, min}}}$
@@
For each $\text{ind}_{i}$, we leverage dynamic programming to further optimize the layer-device scheduling scheme $S$, defining a two-dimensional dynamic programming table $T$, where $T[i][j]$ represents the min-max execution time when assigning the first $i$ layers to the first $j$ devices. The state transition recurrence is given by:
(26) $T[i][j]=\min_{k<i}\left(\max\left(T[k][j-1],\ T_{\text{stage}}(k+1,i,D_{j})\right)\right).$
@@
1) The minimal maximum execution time of the first $k$ layers on first j1j-1 devices, T[k][j1]T[k][j-1].
@@
(27) $T_{\text{stage}}(k+1,\ i,\ D_{j})=\max\left\{\begin{aligned} &T_{\text{comp}}(k+1,\ i,\ D_{j}),\\ &T_{\text{comm}}(D_{j-1},\ D_{j})\end{aligned}\right\}.$
@@
(28) $T_{\text{comp}}(k+1,\ i,\ D_{j})=\sum_{l=k+1}^{i}T_{\text{comp}}(l,\ D_{j}),$
@@
(29) $T_{\text{comm}}(D_{j-1},D_{j})=\begin{cases}\frac{\text{output}_{k}\cdot B_{\mu}}{b_{j-1,j}},&\text{if }D_{j-1}\neq D_{j},\\ 0,&\text{if }D_{j-1}=D_{j}.\end{cases}$
@@
4.4. Theoretical Guarantee of Throughput Improvement under Dataset Variations
While input variations rarely impact throughput in standard inference with fixed model structure and hardware, distributed GNN inference may exhibit performance shifts due to changes in graph structure induced by the dataset. To address this, we derive a theoretical lower bound on throughput improvement under dataset variations, providing formal efficiency guarantees of GA-DPHDS with respect to the indirect impact of datasets.
Problem Setup: We consider a fixed model $\mathcal{M}$ consisting of $L$ layers $\{l_{1},l_{2},\dots,l_{L}\}$ deployed over a set of $K$ devices $\mathcal{D}=\{d_{1},\dots,d_{K}\}$. The micro-batch size $B_{\mu}$, network bandwidths $b_{i-1,i}$, and model architecture are fixed. The only varying factor is the input dataset $D\in\mathcal{D}_{\text{data}}$, which affects the runtime profile of each layer. We define:
  • $t_{j}^{D}$: compute time of layer $l_{j}$ under dataset $D$.
  • $o_{j}^{D}$: output tensor size of layer $l_{j}$.
  • $o_{\max}^{D}=\max_{j}o_{j}^{D}$; $o_{\text{avg}}^{D}=\frac{1}{L}\sum_{j}o_{j}^{D}$: maximum and average output tensor size.
  • $\alpha_{D}:=\frac{o_{\max}^{D}}{o_{\text{avg}}^{D}}$: output tensor imbalance ratio.
  • $b^{\text{up}}_{k}$, $b^{\text{down}}_{k}$: uplink/downlink bandwidth of device $d_{k}$.
  • $\beta:=\frac{\min(b^{\text{good}})}{\min(b^{\text{bad}})}$: link bandwidth asymmetry ratio.
  • $\delta$: the small residual imbalance that remains after scheduling, due to layer indivisibility and hardware heterogeneity.
Pipeline Stage Time Model: For any stage $S_{k}$ allocated to device $d_{k}$, the execution time is:
(30) $T_{k}^{D}=\max\left(\sum_{j\in S_{k}}t_{j}^{D},\quad\frac{o_{j_{\text{last}}}^{D}\cdot B_{\mu}}{\min(b^{\text{up}}_{k-1},b^{\text{down}}_{k})}\right),$
where $j_{\text{last}}$ is the final layer in $S_{k}$.
Throughput Definition: Let $(O,S)$ be a layer ordering and device assignment:
(31) $T_{\text{max}}^{D}(O,S)=\max_{k}T_{k}^{D},\text{Throughput}(O,S;D)=\frac{1}{T_{\text{max}}^{D}(O,S)}.$
We compare two scheduling strategies:
  • Baseline: $(O_{0},S_{0})$ (e.g., equal layer partitioning)
  • GA-DPHDS: $(O^{*},S^{*})$ (optimized via genetic search)
Define the throughput improvement ratio:
(32) $\gamma_{D}:=\frac{\text{Throughput}(O^{*},S^{*};D)}{\text{Throughput}(O_{0},S_{0};D)}=\frac{T_{\text{max}}^{D}(O_{0},S_{0})}{T_{\text{max}}^{D}(O^{*},S^{*})}.$
Step 1: Baseline Bottleneck Estimation In baseline scheduling, stage boundaries are fixed and may split at layers with large output tensors. If such a split falls on a poor bandwidth link, the worst-case stage time becomes:
(33) $T_{\text{baseline}}(D)\geq\frac{o_{\max}^{D}\cdot B_{\mu}}{\min(b^{\text{bad}})},$
which can dominate stage runtime if the split is poorly chosen.
Step 2: GA-DPHDS Improvement GA-DPHDS explicitly searches to avoid such unfavorable splits. Assuming it avoids high-output layers and selects fast links, its worst stage time is upper bounded by:
(34) $T_{\text{GA}}(D)\leq\frac{o_{\text{avg}}^{D}\cdot B_{\mu}}{\min(b^{\text{good}})}+\delta.$
Step 3: Lower Bound on Improvement Ratio Define $\alpha_{D}:=\frac{o_{\max}^{D}}{o_{\text{avg}}^{D}}$ and $\beta:=\frac{\min(b^{\text{good}})}{\min(b^{\text{bad}})}$, then:
(35) $\gamma_{D}\geq\frac{o_{\max}^{D}/\min(b^{\text{bad}})}{o_{\text{avg}}^{D}/\min(b^{\text{good}})+\delta/B_{\mu}}=\frac{\alpha_{D}\cdot\beta}{1+\frac{\delta}{B_{\mu}\cdot o_{\text{avg}}^{D}/\min(b^{\text{good}})}}.$
This gives an explicit lower bound on the throughput improvement, driven by $\alpha_{D}$, $\beta$, and $\delta$.
Theorem 4.1 (Communication-Aware Throughput Gain).
Let model $M$, device cluster $\mathcal{D}$, and micro-batch size $B_{\mu}$ be fixed. For any graph-structured traffic data (dataset) $D$ where $\alpha_{D}$ is bounded, $\beta>1$, and $\delta$ is small enough, then GA-DPHDS scheduling achieves throughput improvement:
(36) $\gamma_{D}\geq\frac{\alpha_{D}\cdot\beta}{1+\varepsilon},\quad\text{where }\varepsilon=\frac{\delta}{B_{\mu}\cdot o_{\text{avg}}^{D}/\min(b^{\text{good}})}\ll 1.$
Practical Significance: Empirical Estimation In our experimental setup shown in Table 1, device bandwidths range from 1.2–3.2 Gbps, i.e., 150–400 MB/s. A realistic range for links yields:
$\min(b^{\text{bad}})=150\text{MB/s},\ \min(b^{\text{good}})=200\text{MB/s}\Rightarrow\beta\approx 1.33.$
In our experiments, $o_{\text{avg}}^{D}=1.11$MB, $o_{\max}^{D}=3.60$MB $\Rightarrow\alpha_{D}=3.24$, and $B_{\mu}=16$.
Estimated communication time under GA:
(37) $\frac{o_{\text{avg}}^{D}}{\min(b^{\text{good}})}=\frac{1.11}{200}=0.00555\text{ s}.$
Assume $\delta=10$ms =0.005=0.005 (converted to seconds), then:
(38) $\gamma_{D}\gtrapprox\frac{3.24\cdot 1.33}{1+0.01/0.00555}=\frac{4.309}{2.802}\approx 1.54.$
Conclusion: GA-DPHDS can theoretically yield at least 1.54× throughput improvement under realistic profile variation and bandwidth heterogeneity, which aligns well with the experimental gain 2.19× of the cluster 4 reported in Section 5.3.1.
This confirms that the theoretical advantage of GA-DPHDS remains practically significant under real hardware constraints.
5. Experiments
The experiments section aims to: (1) Validate the performance of MultiGran-STGCN. (2) Demonstrate the necessity of GA-DPHDS against homogeneous counterparts. (3) Explore the impact of device network bandwidth on overall inference throughput and identify system bottlenecks. (4) Evaluate the effectiveness of our optimization approach.
Table 1. Heterogeneous device table